OpenStudy (anonymous):

if dorquisha had 3 apples and she gave half an apple how many apples does dorquish ahave ? help

OpenStudy (phi):

any idea ?

OpenStudy (phi):

if dorquisha had 1 apple and gave away half of an apple, how many apples does she have ?

OpenStudy (anonymous):

half .. i think

OpenStudy (anonymous):

that dorquisha girl is fat

OpenStudy (phi):

so if she started with 3 apples, and gave away half of an apple, how many apples does she have ?

OpenStudy (anonymous):

2 and a half :D ..?

OpenStudy (phi):

say she put 2 apples in her pocket and kept 1 apple in her hand. then she gives away half of the apple in her hand. She will have half of an apple left in her hand, and 2 more apples still in her pocket. how many apples is that ?

OpenStudy (phi):

yes 2 and 1/2 apples

OpenStudy (phi):

now learn how to do it using fractions: \[ 3 - \frac{1}{2} \] change 3 to 6/2 \[ \frac{6}{2} - \frac{1}{2} \] subtract the tops to get \[ \frac{5}{2} \] make into a mixed number by dividing 2 into 5 to get 2 remainder 1 write that as \[ 2 \frac{1}{2} \]

OpenStudy (anonymous):

oh okay thanks i got it :D